# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 6.2. Fourth moment bounds for $\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)$
(6.3) $\begin{split}{\mathcal{M}}_{N,\varepsilon}^{\varphi,\psi}:&={\mathbbm{E}}\Big{[}\big{(}\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)-{\mathbbm{E}}\big{[}\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)\big{]}\big{)}^{4}\Big{]}=\varepsilon^{2}{\mathbbm{E}}\bigg{[}\bigg{(}\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\quad\quad\quad\quad\quad\varphi_{\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{1})\times\Big{\{}\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j})\Big{\}}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{\varepsilon}(\mathrm{b})\bigg{)}^{4}\bigg{]}.\end{split}$
@@
(6.4) $\begin{split}{\mathcal{M}}_{N,\varepsilon}^{\varphi,\psi}\leq C\|\psi\|_{\infty}^{4}\varepsilon^{3}&\sum\limits_{\mathbf{n}=1}^{2/\varepsilon}{\mathbbm{E}}\bigg{[}\bigg{(}\sum\limits_{r=1}^{\infty}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}\atop|\mathrm{a}|\leq M_{\varepsilon}\sqrt{\mathrm{i}_{1}},|\mathrm{b}|\leq M_{\varepsilon}\sqrt{\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}_{r}^{\prime}}}\\ &\varphi_{\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{1})\times\Big{\{}\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j})\Big{\}}g_{(\mathbf{n}-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\mathbbm{1}_{B_{\varepsilon}}(\mathrm{b})\bigg{)}^{4}\bigg{]}.\end{split}$
@@
(1) Sum over the set $\bigcup_{j=1}^{4}\{\mathrm{i}_{i}^{j}\}:=\{n_{1},\cdots,n_{r}\}$.
@@
Recall that for $\vec{\mathrm{i}}\neq\vec{\mathrm{j}}$, $\Theta_{N,\varepsilon}(\vec{\mathrm{i}})$ and $\Theta_{N,\varepsilon}(\vec{\mathrm{j}})$ are uncorrelated. Since they also have zero means, a summand in (1)-(2) above is nonzero if and only if for any $n_{i}$, we must have $|J_{n_{i}}|\geq 2$.
We also have the following constraints on $(n_{1},\cdots,n_{r})$ and $J_{1},\cdots,J_{r}$ by $(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}$:
(a) $K_{\varepsilon}\leq n_{1}<n_{2}<\cdots<n_{r}\leq\frac{1}{\varepsilon}-K_{\varepsilon}$ with $K_{\varepsilon}=(\log\frac{1}{\varepsilon})^{6}$ (recall (2.38)).
@@
Starting from the constraint (b) above, we can further group the blocks ${\mathcal{D}}_{1},\cdots,{\mathcal{D}}_{m}$. Suppose that for consecutive ${\mathcal{D}}_{i_{1}},{\mathcal{D}}_{i_{1}+1},\cdots,{\mathcal{D}}_{i_{2}}$ ($i_{1}$ may equal to $i_{2}$, that is, we have a single $D_{i_{1}}$), and for any $n_{j}\in\bigcup_{i=i_{1}}^{i_{2}}{\mathcal{D}}_{i}$, we always have that $J_{n_{j}}=\{k,l\}$ with $k\neq l$. Then we only record $s:=\min{\mathcal{D}}_{i_{1}}$ and $t:=\max{\mathcal{D}}_{i_{2}}$ and identify the integer sequence $\bigcup_{i=i_{1}}^{i_{2}}{\mathcal{D}}_{i}$ by $[s,t]$ (similar to the construction of (4.11)). For other (single) ${\mathcal{D}}_{i}$, we also denote $s_{i}=\min{\mathcal{D}}_{i}$ and $t_{i}=\max{\mathcal{D}}_{i}$. Then we can further group ${\mathcal{D}}_{1},\cdots,{\mathcal{D}}_{m}$ according to intervals ${\mathcal{I}}_{i}=[s_{i},t_{i}]\cap{\mathbbm{N}}$, $1\leq i\leq k$. We can now rearrange the summing strategy (1) and (2) as follows:
@@
(2’) Suppose that ${\mathcal{I}}_{i}=[s_{i},t_{i}]$ represents some blocks $\bigcup_{i=i_{1}}^{i_{2}}{\mathcal{D}}_{i}$ with $\min{\mathcal{D}}_{i_{1}}=s_{i}$ and $\max{\mathcal{D}}_{i_{2}}=t_{i}$. Recall the indices set $J_{n_{i}}$ introduced in (2) above. We write ${\mathcal{J}}_{i}=\bigcup_{n_{i}}J_{n_{i}}$, where $n_{i}$ runs over all renewal times in $\bigcup_{i=i_{1}}^{i_{2}}{\mathcal{D}}_{i}$, that is, ${\mathcal{J}}_{i}$ records all indices of the mesoscopic renewal processes that renew at least once at some time $n_{i}\in\bigcup_{i=i_{1}}^{i_{2}}{\mathcal{D}}_{i}$. If $|{\mathcal{J}}_{i}|=2$, then we call the related ${\mathcal{I}}_{i}$ a group of type U, since it is a counterpart of (4.11). If $|{\mathcal{J}}_{i}|\geq 3$, then we call the related ${\mathcal{I}}_{i}$ a a group of type V and it is a counterpart of $\zeta_{n}$ in the expansion of ${\mathcal{Z}}_{N}^{\beta_{N}}(\varphi,\psi)$. Note that as we discussed above, two consecutive U blocks must have different indices set ${\mathcal{J}}_{i}$ and ${\mathcal{J}}_{i+1}$, while a V block must have length $|{\mathcal{I}}_{i}|\leq 4$. Finally, for each ${\mathcal{I}}_{i}=[s_{i},t_{i}]$, we sum over the indices set ${\mathcal{J}}_{i}$.
@@
(A) we unbind $X_{d,f}$ and $X_{d^{\prime},f^{\prime}}$ in the definition of $\Theta_{N,\varepsilon}(\vec{\mathrm{i}})$ (see (2.43)), when $|\vec{\mathrm{i}}|>1$ and $\mathrm{i},\mathrm{i}^{\prime}$ belong to two different groups ${\mathcal{I}}_{i},{\mathcal{I}}_{i^{\prime}}$, so that we have two coarse-grained disorder $\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{1})$ and $\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{2})$ with $\vec{\mathrm{i}}_{1}=(\mathrm{i},\mathrm{i})$ and $\vec{\mathrm{i}}_{2}=(\mathrm{i}^{\prime},\mathrm{i}^{\prime})$.
@@
For step (A), note that since ${\mathcal{I}}_{i}$ and ${\mathcal{I}}_{i^{\prime}}$ are different groups, we have that $s_{i^{\prime}}-t_{i}\geq 2$ Hence, $d^{\prime}-f\geq(s_{i^{\prime}}-t_{i}-1)\varepsilon N\geq\varepsilon N$. Moreover, If ${\mathcal{I}}_{i}$ is a U group, then $\mathrm{i}=t_{i}$. Otherwise, if $\mathrm{i}<t_{i}$, then $\mathrm{i}<\mathrm{i}^{\prime}\leq t_{i}$ and thus $\mathrm{i}^{\prime}$ cannot belong to another group ${\mathcal{I}}_{i^{\prime}}$. If ${\mathcal{I}}_{i}$ is a V group, then $\mathrm{i}\geq s_{i}\geq t_{i}-3$, since the length of a V group can at most be $4$. Similarly, if ${\mathcal{I}}_{i^{\prime}}$ is a U group, then $\mathrm{i}^{\prime}=s_{i^{\prime}}$ and if ${\mathcal{I}}_{i^{\prime}}$ is a V group, then $\mathrm{i}^{\prime}\leq t_{i^{\prime}}\leq s_{i^{\prime}}+3$. In all cases, we have that $d^{\prime}-f\leq(t_{i^{\prime}}+1-s_{i})\leq(s_{i^{\prime}}-t_{i}+7)\varepsilon N$. By local limit theorem, we have that uniformly in all $f\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i})$ and $d^{\prime}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}^{\prime})$, there exists some $c\in(0,+\infty)$, such that
@@
Note that we are considering the moment bounds and we will bound every moment of $\Theta_{N,\varepsilon}(\vec{\mathrm{i}})$ by its absolute value. Hence, we can upper bound any moment of $\Theta_{N,\varepsilon}((\mathrm{i},\mathrm{i}^{\prime}))$ above by the related moment of $C^{\prime}\Theta_{N,\varepsilon}((\mathrm{i},\mathrm{i}))g_{(s_{i^{\prime}}-t_{i})}(0)\Theta_{N,\varepsilon}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))$. We will see that the constant $C^{\prime}$ is not important.
@@
Now, by Chapman-Kolmogorov equation, for any $t:=u_{0}<u_{1}<\cdots<u_{k}=:s$, we have that
$\begin{split}g_{s-t}(0)&=\idotsint\limits_{x_{1},\cdots,x_{k-1}\in{\mathbbm{R}}}g_{u_{1}-u_{0}}(x_{1})\prod\limits_{i=2}^{k-1}g_{(u_{i}-u_{i-1})}(x_{i}-x_{i-1})g_{(u_{k}-u_{k-1})}(x_{k-1}){\mathrm{d}}x_{1}\cdots{\mathrm{d}}x_{k-1}\\ &\leq C^{k}\sum\limits_{z_{1},\cdots,z_{z-1}\in{\mathbbm{Z}}}g_{u_{1}-u_{0}}(z_{1})\prod\limits_{i=2}^{k-1}g_{(u_{i}-u_{i-1})}(z_{i}-z_{i-1})g_{(u_{k}-u_{k-1})}(z_{k-1}),\end{split}$
for some $C$ uniformly. By this decomposition, we introduce a “virtual spatial position”, so that when a mesoscopic renewal process does not renew, we can allocate a “position” to it. Then, we can treat the summation similar to the proof of Theorem 4.1.
@@
(6.5) $\overline{U}_{N,\varepsilon}^{(\rm cg)}(t-s):=\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}\atop\mathrm{i}_{1}=s,\mathrm{i}^{\prime}_{r}=t}{\mathbbm{E}}\big{[}(\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{1}))^{2}\big{]}\prod\limits_{j=2}^{r}g^{2}_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0){\mathbbm{E}}\big{[}(\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j}))^{2}\big{]}.$
@@
(6.6) $\overline{U}_{\infty,\varepsilon}^{\rm(cg)}(t-s):=\lim\limits_{N\to\infty}\overline{U}_{N,\varepsilon}^{(\rm cg)}(t-s).$
@@
(6.7) $V_{N,\varepsilon}^{\rm(cg)}({\mathcal{J}}):={\mathbbm{E}}\Big{[}\prod_{j\in{\mathcal{J}}}\Theta_{N,\varepsilon}((\mathrm{i}_{j},\mathrm{i}^{\prime}_{j}))\Big{]}.$
By Hólder’s inequality and Lemma 6.2, we have that
(6.8) $\limsup_{N\to\infty}V_{N,\varepsilon}^{\rm(cg)}({\mathcal{J}})\leq\limsup_{N\to\infty}\prod\limits_{j\in{\mathcal{J}}}{\mathbbm{E}}\big{[}\Theta_{N,\varepsilon}((\mathrm{i}_{j},\mathrm{i}^{\prime}_{j}))^{4}\big{]}^{\frac{1}{4}}:=V_{\infty,\varepsilon}^{\rm(cg)}({\mathcal{J}})\leq\frac{C}{(\log\frac{1}{\varepsilon})^{\frac{3}{4}}}.$
@@
(3”) We sum over $r\geq 1$ (note that the $r=0$ refers to the mean, which has been subtracted).
@@
(6.9) ${\mathcal{Q}}_{t}(\boldsymbol{x},\boldsymbol{x}^{\prime}):=\prod\limits_{i=1}^{4}g_{t}(x^{\prime}_{i}-x_{i}),\quad{\mathcal{Q}}_{t}(\varphi,\boldsymbol{x}^{\prime}):=\prod\limits_{i=1}^{4}\Big{(}\sum\limits_{y_{i}\in{\mathbbm{Z}}}\varphi(y_{i})g_{t}(x^{\prime}_{i})\Big{)},\quad{\mathcal{Q}}_{t}(\boldsymbol{x},\psi):=\prod\limits_{i=1}^{4}\Big{(}\sum\limits_{y^{\prime}_{i}\in{\mathbbm{Z}}}g_{t}(x_{i})\psi(y^{\prime}_{i})\Big{)}.$

### v3

@@ 6.1 Moment bounds for the coarse-grained variable $\Theta_{N,\varepsilon}$
$\begin{split}{\mathcal{M}}_{N,\varepsilon}^{\varphi,\psi}&:={\mathbbm{E}}\Big[\big(\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)-{\mathbbm{E}}\big[\mathscr{L}_{\varepsilon}^{\text{\rm(cg)}}(\varphi,\psi)\big]\big)^{4}\Big]=\varepsilon^{2}{\mathbbm{E}}\bigg[\bigg(\sum\limits_{r=1}^{(\log\frac{1}{\varepsilon})^{2}}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}}\\ &\quad\quad\quad\varphi_{\varepsilon}(\mathrm{a})g_{\mathrm{i}_{1}}(\mathrm{a})\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{1})\times\Big\{\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j})\Big\}g_{(\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\psi_{\varepsilon}(\mathrm{b})\bigg)^{4}\bigg].\end{split}$ (6.3)
@@
$\begin{split}{\mathcal{M}}_{N,\varepsilon}^{\varphi,\psi}\leq C&\|\psi\|_{\infty}^{4}\varepsilon^{3}\sum\limits_{\mathbf{n}=1}^{2/\varepsilon}{\mathbbm{E}}\bigg[\bigg(\sum\limits_{r=1}^{\infty}\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}}\sum\limits_{\mathrm{a},\mathrm{b}\in{\mathbbm{Z}}\atop|\mathrm{a}|\leq M_{\varepsilon}\sqrt{\mathrm{i}_{1}},|\mathrm{b}|\leq M_{\varepsilon}\sqrt{\lfloor\frac{1}{\varepsilon}\rfloor-\mathrm{i}_{r}^{\prime}}}\\ \varphi_{\varepsilon}(\mathrm{a})&g_{\mathrm{i}_{1}}(\mathrm{a})\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{1})\times\Big\{\prod_{j=2}^{r}g_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0)\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j})\Big\}g_{(\mathbf{n}-\mathrm{i}^{\prime}_{r})}(\mathrm{b})\mathbbm{1}_{B_{\varepsilon}}(\mathrm{b})\bigg)^{4}\bigg].\end{split}$ (6.4)
@@
(1) Sum over the set $\bigcup_{j=1}^{4}\bigcup_{i=1}^{r_{j}}\{\mathrm{i}_{i}^{j}\}:=\{n_{1},\cdots,n_{r}\}$ (all renewal times).
@@
[…]
(a) $K_{\varepsilon}\leq n_{1}<n_{2}<\cdots<n_{r}\leq\frac{1}{\varepsilon}-K_{\varepsilon}$ (recall $K_{\varepsilon}$ from (2.38)).
@@
Starting from the constraint (b) above, we can further group the blocks ${\mathcal{D}}_{1},\cdots,{\mathcal{D}}_{m}$. If for consecutive ${\mathcal{D}}_{i_{1}},{\mathcal{D}}_{i_{1}+1},\cdots,{\mathcal{D}}_{i_{2}}$ ($i_{1}$ may equal to $i_{2}$, that is, we have a single $D_{i_{1}}$), and for any $n_{j}\in\bigcup_{i=i_{1}}^{i_{2}}{\mathcal{D}}_{i}$, we always have that $J_{n_{j}}=\{k,l\}$ with $k\neq l$, then we only record $s:=\min{\mathcal{D}}_{i_{1}}$ and $t:=\max{\mathcal{D}}_{i_{2}}$ and identify the integer sequence $\bigcup_{i=i_{1}}^{i_{2}}{\mathcal{D}}_{i}$ by $[s,t]$ (similar to the construction of (4.11)). For other (single) ${\mathcal{D}}_{i}$, we also denote $s_{i}=\min{\mathcal{D}}_{i}$ and $t_{i}=\max{\mathcal{D}}_{i}$. Then we can further group ${\mathcal{D}}_{1},\cdots,{\mathcal{D}}_{m}$ according to intervals ${\mathcal{I}}_{i}=[s_{i},t_{i}]\cap{\mathbbm{N}}$, $1\leq i\leq k$. We can now rearrange the summing strategy (1) and (2) as follows:
@@
(2’) Suppose that ${\mathcal{I}}_{i}=[s_{i},t_{i}]$ represents some blocks $\bigcup_{i=i_{1}}^{i_{2}}{\mathcal{D}}_{i}$ with $\min{\mathcal{D}}_{i_{1}}=s_{i}$ and $\max{\mathcal{D}}_{i_{2}}=t_{i}$. Recall the indices set $J_{n_{i}}$ introduced in (2) above. We write ${\mathcal{J}}_{i}=\bigcup_{n_{i}}J_{n_{i}}$, where $n_{i}$ runs over all renewal times in $\bigcup_{i=i_{1}}^{i_{2}}{\mathcal{D}}_{i}$, that is, ${\mathcal{J}}_{i}$ records all indices of the mesoscopic renewal processes that renew at least once at some time $n_{i}\in\bigcup_{i=i_{1}}^{i_{2}}{\mathcal{D}}_{i}$. If $|{\mathcal{J}}_{i}|=2$, then we call the related ${\mathcal{I}}_{i}$ a group of type U, since it is a counterpart of (4.11). If $|{\mathcal{J}}_{i}|\geq 3$, then we call the related ${\mathcal{I}}_{i}$ a group of type V and it is a counterpart of $\zeta_{n}$ in the expansion of ${\mathcal{Z}}_{N}^{\beta_{N}}(\varphi,\psi)$. Note that as we discussed above, two consecutive U blocks must have different indices set ${\mathcal{J}}_{i}$ and ${\mathcal{J}}_{i+1}$, while a V block must have length $|{\mathcal{I}}_{i}|\leq 4$. Finally, for each ${\mathcal{I}}_{i}=[s_{i},t_{i}]$, we sum over the indices set ${\mathcal{J}}_{i}$.
@@
(A) We unbind $X_{d,f}$ and $X_{d^{\prime},f^{\prime}}$ in the definition of $\Theta_{N,\varepsilon}(\vec{\mathrm{i}})$ (see (2.43)) when $|\vec{\mathrm{i}}|>1$ and $\mathrm{i},\mathrm{i}^{\prime}$ belong to two different groups ${\mathcal{I}}_{i},{\mathcal{I}}_{i^{\prime}}$, so that we have two coarse-grained disorder $\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{1})$ and $\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{2})$ with $\vec{\mathrm{i}}_{1}=(\mathrm{i},\mathrm{i})$ and $\vec{\mathrm{i}}_{2}=(\mathrm{i}^{\prime},\mathrm{i}^{\prime})$.
@@
For step (A), note that since ${\mathcal{I}}_{i}$ and ${\mathcal{I}}_{i^{\prime}}$ are different groups, we have that $s_{i^{\prime}}-t_{i}\geq 2$ . Hence, $d^{\prime}-f\geq(s_{i^{\prime}}-t_{i}-1)\varepsilon N\geq\varepsilon N$. Moreover, If ${\mathcal{I}}_{i}$ is a U group, then $\mathrm{i}=t_{i}$. Otherwise, if $\mathrm{i}<t_{i}$, then $\mathrm{i}<\mathrm{i}^{\prime}\leq t_{i}$ since the indices of renewal processes in a U group cannot change, and thus $\mathrm{i}^{\prime}$ cannot belong to another group ${\mathcal{I}}_{i^{\prime}}$. If ${\mathcal{I}}_{i}$ is a V group, then $\mathrm{i}\geq s_{i}\geq t_{i}-3$, since the length of a V group can at most be $4$. Similarly, if ${\mathcal{I}}_{i^{\prime}}$ is a U group, then $\mathrm{i}^{\prime}=s_{i^{\prime}}$ and if ${\mathcal{I}}_{i^{\prime}}$ is a V group, then $\mathrm{i}^{\prime}\leq t_{i^{\prime}}\leq s_{i^{\prime}}+3$. In all cases, we have that $d^{\prime}-f\leq(t_{i^{\prime}}+1-s_{i})\leq(s_{i^{\prime}}-t_{i}+7)\varepsilon N$. By local limit theorem, we have that uniformly in all $f\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i})$ and $d^{\prime}\in{\mathcal{T}}_{N,\varepsilon}(\mathrm{i}^{\prime})$, there exists some $c\in(0,+\infty)$, such that
@@
Note that we are considering the moment bounds and we will bound every moment of $\Theta_{N,\varepsilon}(\vec{\mathrm{i}})$ by its absolute value, which can be done by bounding the absolute value of moments of all $\zeta_{n}$’s (see Theorem 4.1). Hence, we can upper bound any moment of $\Theta_{N,\varepsilon}((\mathrm{i},\mathrm{i}^{\prime}))$ above by the related moment of $C^{\prime}\Theta_{N,\varepsilon}((\mathrm{i},\mathrm{i}))g_{(s_{i^{\prime}}-t_{i})}(0)\Theta_{N,\varepsilon}((\mathrm{i}^{\prime},\mathrm{i}^{\prime}))$. We will see that the constant $C^{\prime}$ is not important.
@@
By Chapman-Kolmogorov equation, for any $t:=u_{0}<u_{1}<\cdots<u_{k}=:s$, we have that
$\begin{split}g_{s-t}(0)&=\idotsint\limits_{x_{1},\cdots,x_{k-1}\in{\mathbbm{R}}}g_{u_{1}-u_{0}}(x_{1})\prod\limits_{i=2}^{k-1}g_{(u_{i}-u_{i-1})}(x_{i}-x_{i-1})g_{(u_{k}-u_{k-1})}(-x_{k-1}){\mathrm{d}}x_{1}\cdots{\mathrm{d}}x_{k-1}\\ &\leq C^{k}\sum\limits_{z_{1},\cdots,z_{z-1}\in{\mathbbm{Z}}}g_{u_{1}-u_{0}}(z_{1})\prod\limits_{i=2}^{k-1}g_{(u_{i}-u_{i-1})}(z_{i}-z_{i-1})g_{(u_{k}-u_{k-1})}(z_{k-1}),\end{split}$
for some $C$ uniformly. By this decomposition, we introduce “virtual spatial positions”, so that when a mesoscopic renewal process does not renew, we can allocate a “position” to it. Then, we can treat the summation similar to the proof of Theorem 4.1.
@@
$\overline{U}_{N,\varepsilon}^{(\rm cg)}(t-s):=\sum\limits_{(\vec{\mathrm{i}}_{1},\cdots,\vec{\mathrm{i}}_{r})\in\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}\atop\mathrm{i}_{1}=s,\mathrm{i}^{\prime}_{r}=t}{\mathbbm{E}}\big[(\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{1}))^{2}\big]\prod\limits_{j=2}^{r}g^{2}_{(\mathrm{i}_{j}-\mathrm{i}^{\prime}_{j-1})}(0){\mathbbm{E}}\big[(\Theta_{N,\varepsilon}(\vec{\mathrm{i}}_{j}))^{2}\big].$ (6.5)
@@
$\overline{U}_{\infty,\varepsilon}^{\rm(cg)}(t-s):=\lim\limits_{N\to\infty}\overline{U}_{N,\varepsilon}^{(\rm cg)}(t-s).$ (6.6)
@@
$V_{N,\varepsilon}^{\rm(cg)}({\mathcal{J}}):={\mathbbm{E}}\Big[\prod_{j\in{\mathcal{J}}}\Theta_{N,\varepsilon}((\mathrm{i}_{j},\mathrm{i}^{\prime}_{j}))\Big].$ (6.7)
By Hölder’s inequality and Lemma 6.2, we have that
$\limsup_{N\to\infty}V_{N,\varepsilon}^{\rm(cg)}({\mathcal{J}})\leq\limsup_{N\to\infty}\prod\limits_{j\in{\mathcal{J}}}{\mathbbm{E}}\big[\Theta_{N,\varepsilon}((\mathrm{i}_{j},\mathrm{i}^{\prime}_{j}))^{4}\big]^{\frac{1}{4}}:=V_{\infty,\varepsilon}^{\rm(cg)}({\mathcal{J}})\leq\frac{C}{(\log\frac{1}{\varepsilon})^{\frac{3}{4}}}.$ (6.8)
@@
(3”) We sum over $r\geq 1$ (note that $r=0$ refers to the mean, which has been subtracted).
@@
${\mathcal{Q}}_{t}(\boldsymbol{x},\boldsymbol{x}^{\prime})$ $:=\prod\limits_{i=1}^{4}g_{t}(x^{\prime}_{i}-x_{i}),$ (6.9)
${\mathcal{Q}}_{t}(\varphi,\boldsymbol{x}^{\prime})$ $:=\prod\limits_{i=1}^{4}\Big(\sum\limits_{y_{i}\in{\mathbbm{Z}}}\varphi(y_{i})g_{t}(x^{\prime}_{i})\Big),$ (6.10)
${\mathcal{Q}}_{t}(\boldsymbol{x},\psi)$ $:=\prod\limits_{i=1}^{4}\Big(\sum\limits_{y^{\prime}_{i}\in{\mathbbm{Z}}}g_{t}(x_{i})\psi(y^{\prime}_{i})\Big).$ (6.11)